Sucrose (C12H22O11, table sugar) is oxidized in the body

by O2 via a complex set of reactions that ultimately produces

CO2(g) and H2O(g) and releases 5.64 103 kJ/mol sucrose.

(a) Write a balanced thermochemical equation for this reaction.

(b) How much heat is released per gram of sucrose oxidized?


1
Expert's answer
2021-11-25T01:28:52-0500

(a). Balanced thermochemical equation for this reaction is :

C12H22O11(s) + 12 O2(g) --------------> 12 CO2(g) + 11 H2(l) ,  ΔHrxn = - 5.16 × 103 kJ

(b).

For the combustion of 1 mole of C12H22O11 , heat released = 5.16 x 103 KJ

Now, Mass of 1 mole of C12H22O11 = Number of moles x Gram molar mass = 1 mol x 342.3 g/mol = 342.3 g

Now, for the combustion of 342.3 g i.e. 1 mole of C12H22O11 ,heat released = 5.16 x 103 KJ

and

For the combustion of 1g of C12H22O11 , heat released = 5.16 x 103 KJ x 1g / 342.3 g = 15.07 KJ

Hence, Energy change per gram of sucrose for this combustion reaction = -15.07 KJ/g